On a particular day, the wind added 5 miles per hour to Jaime's rate when she was rowing with the wind and subtracted 5 miles per hour from her rate on her return trip. Jaime found that in the same amount of time she could row 75 miles with the wind, she could go only 45 miles against the wind.What is her normal rowing speed with no wind?
Answer by macston(5194)

.
n=normal speed(no wind)
.
time with wind:
t(with)=75/(n+5mph)
.
time against wind:
t(against)=45/(n-5mph)
.
Since t(with)=t(against):
75/(n+5mph)=45/(n-5mph) Cross multiply.
75(n-5mph)=45(n+5mph)
75n-375=45n+225 Subtract 45n from each side.
30n-375=225 Add 375 to each side.
30n=600 Divide each side by 30.
n=20 ANSWER: Jaime's normal rowing speed with no wind is 20 mph.
.
CHECK:
n=20 mph
75mi/(n+5mph)=45mi/(n-5mph)
75mi/(20mph+5mph)=45mi/(20mph-5mph)
75mi/25mph=45mi/15mph
3 hrs=3 hrs